One end of a U-tube that contains water is open to the atmosphere but the other end is closed. When some air is trapped in the closed end, the difference in the height of the water, h, is 14 .0 cm. Assume the outside air pressure is 0.9600 atm.

A) What is the gauge pressure of the trapped air?

B) What is the absolute pressure of the trapped air?
Answer #1

We have

p abs = Po+ pgh = Patm +pgh

P= 1.013* 10^5 pascal + 1000kg/ m^3*9.8 m/sec^2)*14/100 mts
= 13.89* 10^5 Pascal

paguge = p - patm

        13.89 * 10^5 pascal - 1.013*10^5 pascal
   
             = 12.877*10^5 pascal
               = 12.877*0.96/1.103*10^5 atm
                   = 12.20 atm

absolute pressure = pgauge + Patm
                    12.20 +0.96 = 13.16 atm